OpenStudy (anonymous):

Please help me!! I need help on matrices. this is the problem. A= find 3A A= {7 2 9 } {8 6 4 } then I did this A= {6 4} {3 10 5 } {10 5} I crossed multiplied and then I subtracted 30-40= -10 Do I now just times it by 3 because it says find 3A or did I do something wrong? please help.

OpenStudy (unklerhaukus):

(just checking the question) Is A this 3x3 matrix ? A= {7, 2, 9} {8, 6, 4} {3, 10, 5} And you are trying to find 3A ?

jimthompson5910 (jim_thompson5910):

to find 3A all you need to do is multiply EVERY component of the matrix A by 3

OpenStudy (anonymous):

Yes! Then I did {6, 4} I crossed multiplied {10, 5}

OpenStudy (unklerhaukus):

im not sure why you have cross multiplied, and i dont know where that 2x2 matrix has come from? it looks like a mix between a few different questions

jimthompson5910 (jim_thompson5910):

ivonlorena, if you start out with a 3x3 matrix, you should end up with a 3x3 matrix

OpenStudy (anonymous):

Alright so basically I multiply 7x3 then 2x3 then 9x3 and so on? or am I doing something wrong again

OpenStudy (anonymous):

Wait I don't know :(

jimthompson5910 (jim_thompson5910):

you are correct multiply every number in the matrix by 3

jimthompson5910 (jim_thompson5910):

and that's all you need to do to find 3A

OpenStudy (anonymous):

so after I multiply each number by 3 what do I do next?

jimthompson5910 (jim_thompson5910):

you're done after that

OpenStudy (unklerhaukus):

if all you were trying to find is 3A , and you have multiplied all the elements of A by three, then you are done.

OpenStudy (anonymous):

Thank you so much UnkleRhaukus and jim_thompson5910!

OpenStudy (anonymous):

I also need help on this one if you don't mind.. Find A+B where B= { 1, 1, 1} { 1, 1, 1} { 1, 1, 1} do I do the same thing?

jimthompson5910 (jim_thompson5910):

now you are adding the matrices so you add up the corresponding elements

jimthompson5910 (jim_thompson5910):

A= {7, 2, 9} {8, 6, 4} {3, 10, 5} B= { 1, 1, 1} { 1, 1, 1} { 1, 1, 1}

jimthompson5910 (jim_thompson5910):

So in the upper right hand corner for instance, you'll have 9+1 = 10 The answer will be a 3x3 matrix and in the upper right hand corner of the answer matrix will have 10 in it. The other entries are found the same way (by adding up the corresponding elements)

OpenStudy (anonymous):

So it would be? {8, 3, 10} {9, 6, 5} {4, 11, 6}

jimthompson5910 (jim_thompson5910):

everything is good, but the very center it's not 6, it should be 7

OpenStudy (anonymous):

oh woops that was an accident lol

jimthompson5910 (jim_thompson5910):

that's ok

OpenStudy (anonymous):

So in the last one it says Find 2(AC) where c= {-3 -3 -3} {-3 -3 -3} {-3 -3 -3}

jimthompson5910 (jim_thompson5910):

ok first you need to multiply matrix A with matrix C however, multiplication isn't as simple as just multiplying the corresponding elements (like you do with addition). Matrix multiplication is a bit trickier. Are you familiar with it?

OpenStudy (anonymous):

Wait nevermind I guess I'm not familiar with it

OpenStudy (anonymous):

It was something like this but I don't know if this is a different kind of problem

jimthompson5910 (jim_thompson5910):

that's a determinant, which is different

jimthompson5910 (jim_thompson5910):

First off, we have to see if we can even multiply A and C A is a 3x3 matrix C is a 3x3 matrix A has 3 columns, C has 3 rows. Because these '3's match up, we can multiply A and C The final result will be a 3x3 matrix (the 3's are drawn from the first 3 of A and the last 3 of C) ----------------------------- The way this is done is that each element in the final matrix will be found separately. To find the element in row1, column1 of the final matrix, you would multiply the elements of row 1 of A with the corresponding elements of the column 1 in C After you find all these products, you'll add them up to get the result for row1, column1

jimthompson5910 (jim_thompson5910):

this image is a good way of thinking about it http://cnx.org/content/m32159/1.4/matrixMultiply.png

OpenStudy (anonymous):

I don't think I understand I ended up with 729. I think I did it completely wrong

jimthompson5910 (jim_thompson5910):

what are the elements in the first row of matrix A?

OpenStudy (anonymous):

-3 -3 -3?

jimthompson5910 (jim_thompson5910):

that's matrix C

OpenStudy (anonymous):

Omg I get it now I know what I did wrong.

OpenStudy (anonymous):

Okay let me do it again. I'm sorry.

OpenStudy (anonymous):

Are the first elements 7, 2, 9 or 7, 8, 3?

jimthompson5910 (jim_thompson5910):

we look at row 1 of matrix A, so 7,2,9

jimthompson5910 (jim_thompson5910):

column 1 of matrix C is -3, -3, -3

jimthompson5910 (jim_thompson5910):

multiply those corresponding elements together 7(-3) = -21 2(-3) = -6 9(-3) = -27 then add up those products -21 + (-6) + (-27) = -54

jimthompson5910 (jim_thompson5910):

the result of -54 is the element that goes in row1, colum1 of the answer matrix this needs to be done 8 more times (9 times total because have 9 elements in AC)

OpenStudy (anonymous):

8x-3=-24 6x-3=-18 4x-3=-12 all together will be -54. So then I put that next to my other answer.

OpenStudy (anonymous):

In the same row1, column 1?

jimthompson5910 (jim_thompson5910):

notice how you're focusing on row 2 of matrix A if you picked on column 1 of matrix C, then -54 will go in row2, column1 of AC

OpenStudy (anonymous):

But I thought it was supposed to be done 9 more times.

jimthompson5910 (jim_thompson5910):

it is, once for each entry in AC matrix AC is a 3x3 matrix (9 entries total)

jimthompson5910 (jim_thompson5910):

row1, column1 for AC is done by picking on row 1 from A, column 1 from C row2, column1 for AC is done by picking on row 2 from A, column 1 from C etc etc

OpenStudy (anonymous):

So then it would just be -54? all of them?

jimthompson5910 (jim_thompson5910):

Let me check, one moment

jimthompson5910 (jim_thompson5910):

Yes, every result I'm getting is -54 However, it's always a good idea to go through and compute all the entries (so you can get practice)

jimthompson5910 (jim_thompson5910):

anyways, after you get AC, you double everything in AC to get 2(AC)

OpenStudy (anonymous):

So -108?

jimthompson5910 (jim_thompson5910):

every entry will have -108, correct

jimthompson5910 (jim_thompson5910):

so in the end, you'll have a 3x3 matrix filled with nothing but -108

OpenStudy (anonymous):

Thank you very much for your help I understand now :)

jimthompson5910 (jim_thompson5910):

you're welcome, I'm glad you do